OpenStudy (anonymous):

darla is building a new desk. to make sure she had made a square corner she measures 10 ft from the corner along one edge and 6ft from the corner of the other edge. how long should the diagonal be between those two points if the corner is a right angle

OpenStudy (anonymous):

sqrt(100+36) =11.7 sq. units

OpenStudy (anonymous):

round to the thousandth

OpenStudy (anonymous):

Since the corner needs to be square, ie 90 degrees, we can use Pythagorean Theorem. \[c^2 = a^2 + b^2\] where c is the length of the diagonal, and a and b are lengths of the two measured sides. With numbers this becomes\[c^2 = 10^2 + 6^2 \rightarrow c = \sqrt{(100 + 36)} = \sqrt{(136)}\]

OpenStudy (anonymous):

\[\sqrt{10^2+6^2}=\sqrt{136}=11.66190379\] round to the thousandth= 11.662